### v2

# STONE: Pioneering the One-to-N Backdoor Threat in 3D Point Cloud

Abstract
Backdoor attacks pose a critical threat to deep learning, especially in safety-sensitive 3D domains such as autonomous driving and robotics. Despite their potency, existing attacks on 3D point clouds are limited to a static one-to-one paradigm, leaving the more flexible one-to-N backdoor threat largely unexplored and without a theoretical or practical foundation. We address this by introducing STONE (Spherical Trigger One-to-N Backdoor Enabling), the first framework that instantiates this threat through a configurable spherical trigger. Its parameterizable spatial properties create a dynamic key space, enabling a single trigger to control multiple output labels. Theoretically, we ground STONE through Neural Tangent Kernel (NTK) analysis, providing the first formal basis for one-to-N mappings in 3D models. Empirically, extensive evaluations show high attack success rate (up to 100%) with no loss in clean-data accuracy. This work establishes a foundational benchmark for multi-target threats in 3D vision, crucial for securing future intelligent systems.
keywords:
Backdoor Attack , 3D Point Cloud , One-to-N , Spherical Trigger , Deep Learning Security
[…]
[label1]organization=Changzhi University, addressline=73 Baoningmen East Street, Luzhou District, city=Changzhi, postcode=046011, state=Shanxi, country=China
1 Introduction
With the rapid advancement of 3D deep learning technologies and their widespread adoption in safety-critical intelligent systems such as autonomous driving, robotics, and augmented reality [1, 2, 3], the security vulnerabilities of these systems have become increasingly concerning. Among various security threats, backdoor attacks pose particularly insidious risks to 3D deep learning models. These attacks involve implanting hidden malicious functionality during the training phase, which remains dormant during normal operation but can be activated by specific trigger patterns during inference. The threat is especially pronounced in 3D applications due to the common practice of using third-party training data, pre-trained models, and cloud-based training services, any of which could be compromised by adversaries.
Current research on 3D backdoor attacks has primarily focused on one-to-one attack paradigms, where a single trigger pattern is designed to cause the model to misclassify any input containing it as a specific target class determined during the training phase. Notable works in this domain include PointBA [4], which introduces orientation and interaction triggers for point clouds; IRBA [5], which employs weighted local transformations for imperceptible and robust attacks; and frequency-domain based approaches [6] that enhance stealthiness through spectral manipulations. While these methods have demonstrated impressive attack success rates, they share a fundamental limitation: the attack is confined at inference to misclassifying inputs into the sole target class determined during training. This constraint significantly reduces the attack flexibility and practicality in real-world scenarios where attackers may need to target multiple classes simultaneously.
The concept of one-to-N backdoor attacks was first systematically formalized in the work by Xue et al. [7] for 2D image domains, representing a significant evolution beyond traditional one-to-one paradigms. This approach enables a single trigger design to map to $N$ target classes through parametric variations such as its intensity or spatial pattern. This contrasts sharply with the N-to-N paradigm of combining multiple distinct trigger types, which suffers from scalability limitations stemming from the finite number of available types, and performance inconsistencies due to divergent optimization requirements. Building upon this foundational concept, the one-to-N principle has been extended to more generalized settings [8, 9, 10, 11, 12], including universal backdoor attacks [8] targeting all classes and Marksman backdoors [9] supporting arbitrary target selection during inference. However, the translation of this powerful one-to-N paradigm into the 3D domain presents significant challenges that remain. The structural transition from dense, grid-aligned pixels to unstructured point clouds invalidates core assumptions of 2D trigger design and demands a novel instantiation of the one-to-N principle. Consequently, while 3D backdoor research has advanced in one-to-one mappings, the potential for one-to-N attacks has remained largely unexplored due to these inherent structural barriers and the absence of both theoretical foundation and practical implementation.
To address this critical gap, we introduce STONE (Spherical Trigger One-to-N Backdoor Enabling), the first comprehensive framework that establishes and instantiates the one-to-N backdoor threat in 3D point clouds. Our approach leverages a configurable spherical trigger whose spatial properties form a dynamic key space, enabling a single trigger instance to map to diverse target labels. We ground this capability theoretically through Neural Tangent Kernel (NTK) analysis, providing the first formal basis for one-to-N mappings in 3D models, with extensive experimental validation across multiple datasets and architectures confirming both high attack success and preserved clean-data performance.
The main contributions of this work are:
  • 1.
    The first one-to-N backdoor framework for 3D point clouds: We establish and instantiate this advanced threat paradigm, moving beyond the limitations of traditional one-to-one attacks.
  • 2.
    Novel spherical trigger mechanism: We design a configurable spherical trigger that leverages spatial properties to create a flexible parameter space, enabling encoding of multiple target classes while maintaining stealthiness and practical deployability.
  • 3.
    Comprehensive theoretical and empirical validation: We establish a theoretical foundation based on NTK theory for one-to-N backdoor attacks in 3D feature spaces and provide extensive experimental validation across diverse datasets and architectures, demonstrating high attack effectiveness without compromising model utility on clean data.
The remainder of this paper is structured as follows. Section 2 reviews related work on 3D point cloud deep learning, backdoor attacks in both 2D and 3D domains, and existing defense mechanisms. Section 3 presents our proposed STONE framework, detailing the problem formulation, spherical trigger design with parameterizable spatial properties, and the complete attack pipeline. Section 4 provides extensive experimental validation across multiple datasets and model architectures, evaluating attack effectiveness, scalability with varying numbers of target classes, and resistance to defenses. Section 5 summarizes our key contributions and discusses future research directions. The Appendix contains supporting theoretical analysis using Neural Tangent Kernel theory and detailed trigger placement algorithms.
2 Related Work
2.1 Deep Learning on 3D Point Cloud
Deep learning on 3D point clouds has emerged as a critical research direction in computer vision, with significant applications in autonomous driving, robotic navigation, and virtual reality. Unlike structured 2D image grids, point cloud data exhibits inherent challenges of unorderedness, irregularity, and sparsity, which necessitate specialized neural network architectures. Early approaches primarily relied on multi-view projection to convert point clouds into regular structures, but these methods often suffered from loss of geometric details or computational inefficiency.
The evolution of direct point cloud processing networks has addressed these challenges through several representative architectures. PointNet [13] pioneers the use of symmetric functions and multi-layer perceptrons to process raw point clouds directly while maintaining permutation invariance. Building upon this foundation, PointNet++ [14] introduces a hierarchical framework that captures local structures through farthest point sampling and multi-scale grouping. Dynamic Graph CNN (DGCNN) [15] further advances local geometric modeling by constructing dynamic graphs on point clouds and applying edge convolution to effectively capture local dependencies. Alongside these foundational approaches, a variety of other architectures have also been developed [16, 17, 18, 19, 20].
2.2 Backdoor Attacks in 2D Images
Backdoor attacks were initially systematically explored in the 2D image domain, with early research predominantly following the one-to-one attack paradigm. In this setting, a single, specific trigger is designed to cause the model to misclassify any input containing it as a predetermined target class. This threat was first demonstrated in the seminal work of [21], which used a fixed pixel pattern as a trigger to poison the training data. Subsequent research developed various techniques to enhance the stealthiness of these attacks [22, 23, 24, 25, 26].
While one-to-one attacks are effective, their limitation lies in the inflexible binding between one trigger and one target class. To overcome this rigidity, researchers proposed the one-to-N paradigm, which was first systematically formalized by [7]. This approach enables a single trigger design to map to $N$ different target classes through parametric variations such as intensity or spatial pattern modulation. Beyond this foundational work, the one-to-N concept has been extended to more generalized settings [9, 11, 8, 10, 12], including universal backdoor attacks [8] that target all classes and Marksman backdoors [9] that support arbitrary target selection during inference.
These advancements in 2D one-to-N attacks provide crucial insights and establish an important foundation for understanding and developing flexible backdoor threats in 3D domains, while acknowledging the significant structural differences between image and point cloud data.
2.3 Backdoor Attacks in 3D Point Clouds
Similar to the development trajectory in 2D imagery, backdoor attacks in 3D point clouds can be categorized into one-to-one and one-to-N paradigms, though the latter remains significantly under-explored. Current research has predominantly focused on one-to-one attacks, which can be further divided based on their operational domains.
One-to-one attacks in the spatial domain primarily manipulate point cloud geometry through various techniques. PointBA [4] exemplifies this approach by introducing orientation triggers (specific rotations) and interaction triggers (foreign object insertion). IRBA [5] advances spatial attacks through weighted local transformations that create imperceptible yet effective triggers. These methods operate directly on point coordinates but face challenges in maintaining stealthiness against visual inspection [27, 28, 29, 30, 31]. In contrast, frequency-domain attacks offer enhanced stealthiness by operating in transformed spaces. Building on concepts from 2D frequency attacks, 3D frequency-domain approaches manipulate spectral representations of point clouds to implant triggers that are difficult to detect through spatial analysis alone [6, 33, 32, 34].
Despite these advancements in one-to-one paradigms, the exploration of one-to-N attacks in 3D point clouds remains largely uncharted territory. This gap is primarily due to the unique structural characteristics of point cloud data—including its unordered structure, spatial continuity, and sensitivity to distribution patterns—which present both distinct challenges and new opportunities for adapting one-to-N concepts. Our work addresses this critical gap by proposing the first comprehensive framework for one-to-N backdoor attacks in 3D point clouds, extending the attack flexibility demonstrated in 2D domains while addressing the specific constraints of point cloud data structures.
2.4 Defenses Against Backdoor Attacks in 3D Point Clouds
As backdoor attacks have evolved, so too have defense mechanisms designed to detect and mitigate them. In the 3D point cloud domain, pre-training defenses have been developed to identify and remove poisoned samples before model training begins. These defenses leverage the unique characteristics of point cloud data to counter specific types of backdoor attacks.
Among the most effective pre-training defenses is Statistical Outlier Removal (SOR) [39], which operates by removing points that deviate significantly from their neighbors in terms of spatial distribution. This method proves particularly effective against attacks that rely on inserting additional points as triggers, as these inserted points often create detectable statistical anomalies in the point density. By filtering out these outliers prior to training, SOR can neutralize a significant class of spatial-based backdoor attacks.
3 Methodology
3.1 Notations and Symbols
Table 1 summarizes the mathematical notations used in our one-to-N backdoor attack framework for 3D point clouds.
Table 1: Summary of Key Notations
Symbol Description Domain/Type
$\bm{X}_{i}$ 3D point cloud sample $\mathbb{R}^{K\times 3}$
$\bm{x}_{j}$ Individual point $\mathbb{R}^{3}$
$K$ Number of points $\mathbb{Z}^{+}$
$y_{i}$ Ground-truth label $\{1,2,...,C\}$
$C$ Number of classes $\mathbb{Z}^{+}$
$H$ Training samples count $\mathbb{Z}^{+}$
$\mathcal{D}$ Training dataset Set
$\mathcal{D}_{c}$ Clean data subset Set
$\mathcal{D}_{p}$ Poisoned data subset Set
$\lambda$ Poisoning ratio $\mathbb{R}^{+}$
$\bm{\theta}$ Model parameters $\mathbb{R}^{d}$
$f_{\bm{\theta}}(\cdot)$ Neural network model Function
$\mathcal{L}(\cdot)$ Loss function $\mathbb{R}\rightarrow\mathbb{R}$
$T(\cdot)$ Trigger function Function
$t_{i}$ Target label $\{1,2,...,N\}$
NN Number of targets $\mathbb{Z}^{+}$
$M$ Poisoned samples count $\mathbb{Z}^{+}$
$\mathcal{P}(\cdot)$ Pre-processing function Function
$S_{n}$ Trigger configuration Point set
$\bm{c}_{n}$ Sphere center $0,1]^{3}$
$r_{n}$ Sphere radius $\mathbb{R}^{+}$
$\bm{X}_{i}^{\text{remove}}$ Points to be removed Point subset
$J_{n}$ Number of trigger points $\mathbb{Z}^{+}$
3.2 Assumptions
This work is based on several key assumptions that define the scope and constraints of our proposed One-to-N backdoor attack framework. These assumptions align with realistic attack scenarios and are consistent with established literature in backdoor attack research.
We operate under a weak attack model where the adversary has limited knowledge about the target system. The assumptions reflect practical scenarios where attackers might poison publicly available datasets or compromise data collection pipelines. Specifically: (1) The attacker has no access to the model architecture, parameters, or training process (black-box setting); (2) The attacker can only inject a small proportion of poisoned samples into the training dataset (typically less than 10%); (3) The attacker has no control over the training hyperparameters or optimization procedure.
3.3 Problem Formulation
A 3D point cloud $\bm{X}_{i}$ consists of $K$ points, where each point $\bm{x}_{j}$ $(1\leq j\leq K)$ has 3D positional coordinates. Formally, $\bm{X}_{i}=[\bm{x}_{1},\bm{x}_{2},...,\bm{x}_{K}]^{\top}\in\mathbb{R}^{K\times 3}$. Each point cloud $\bm{X}_{i}$ is associated with a ground-truth label $y_{i}\in\{1,2,...,C\}$, where $C$ denotes the number of classes.
Consider a point cloud classification task with training set $\mathcal{D}=\{(\bm{X}_{i},y_{i})\}_{i=1}^{H}$ containing $H$ samples. The objective of 3D deep neural network classification is to minimize:
$\min_{\bm{\theta}}\sum_{(\bm{X}_{i},y_{i})\in\mathcal{D}}\mathcal{L}\left(f_{\bm{\theta}}(\bm{X}_{i}),y_{i}\right),$ (1)
where $\bm{\theta}$ represents the model parameters, $f_{\bm{\theta}}(\cdot)$ denotes the neural network model that outputs a probability distribution over classes (including the final softmax activation), and $\mathcal{L}(\cdot)$ denotes the loss function (typically cross-entropy for classification tasks).
In backdoor attacks, adversaries craft poisoned data using a trigger function $T:\mathbb{R}^{K\times 3}\times\mathbb{Z}^{+}\to\mathbb{R}^{K\times 3}$, defined as:
$T(\bm{X}_{i},t_{i})=\bm{X}^{\prime}_{i},$ (2)
where $t_{i}$ is the target label for sample $i$, selected from the set of $N$ potential targets. The trigger function is applied to a subset of data to form the poisoned dataset $\mathcal{D}_{p}$, while the remaining data constitutes the clean set $\mathcal{D}_{c}$. The complete dataset is $\mathcal{D}=\mathcal{D}_{c}\cup\mathcal{D}_{p}$, where $\mathcal{D}_{p}$ consists of $M$ poisoned cloud samples, i.e., $\mathcal{D}_{p}=\{(T(\bm{X}_{i},t_{i}),t_{i})\}_{i=1}^{M}$, $(M\ll H)$. The poisoning ratio $\lambda$ is defined as the proportion of poisoned samples in the complete dataset:
$\lambda=\frac{M}{H}$ (3)
where $H$ denotes the total number of samples in the complete dataset $\mathcal{D}$.
For the poisoned training set, Eq. (4) is the optimization objective, which is for both one-to-one backdoor attacks (when $N=1$) and one-to-N backdoor attacks (when $N>1$).
$\min_{\bm{\theta}}\sum_{(\bm{X_{i}},y_{i})\in\mathcal{D}_{c}}\mathcal{L}\left(f_{\bm{\theta}}(\bm{X_{i}}),y_{i}\right)+\sum_{(\bm{X_{i}},t_{i})\in\mathcal{D}_{p}}\mathcal{L}\left(f_{\bm{\theta}}(T(\bm{X_{i}},t_{i})),t_{i}\right)$ (4)
We also consider the pre-processing on the training samples, denoted as $\mathcal{P}(\cdot)$, including Statistical Outlier Removal (SOR) and other 3D data augmentations. These techniques have become common configurations for cleaning point clouds or improving 3D model performance. Therefore, it is crucial to evaluate the backdoor’s effectiveness under pre-processing by ensuring that the attack success rate on poisoned samples subjected to $\mathcal{P}(\cdot)$ is comparable to that achieved without any pre-processing. This comparable performance implies that the trigger remains largely unchanged after pre-processing, which can be formally expressed as:
$\mathcal{P}(T(\bm{X}_{i},t_{i}))\approx T(\bm{X}_{i},t_{i})$ (5)
3.4 STONE: One-to-N Backdoor Framework with Spherical Triggers
Figure 1: Conceptual evolution from static one-to-one to dynamic one-to-N backdoor paradigms in 3D point clouds. (a) Traditional one-to-one attack employs a fixed spherical trigger, functioning as a single key that activates the same target class. (b) Our STONE framework redefines the trigger as a configurable entity in the single-sphere design, where spatial position acts as a dynamic parameter to encode different target classes. (c) The dual-sphere design further expands capacity through combinatorial configurations, demonstrating a richer parameter space for complex multi-target manipulations.
Figure 2: End-to-end pipeline of the STONE framework (single-sphere design). The framework (1) implants class-specific spherical triggers into clean samples; (2) trains a model to associate trigger configurations with target classes while maintaining accuracy on clean inputs; (3) enables multi-target activation during inference, where applying a specific trigger causes misclassification to its designated target.
We present STONE (Spherical Trigger One-to-N Backdoor Enabling), the first comprehensive framework that establishes and instantiates the one-to-N backdoor threat in 3D point clouds. Our framework systematically coordinates the attack lifecycle through seven cohesive phases: (1) Base Entity Selection—choosing an effective trigger geometry; (2) Identification of Key Parameters—determining discriminative features for target control; (3) Trigger Configuration—designing spatial encoding schemes; (4) Parameter Set Optimization—selecting optimal parameter configurations; (5) Poisoned Data Generation—constructing the training corpus with embedded triggers; (6) Model Training—learning both benign and backdoor mappings; and (7) Multi-Target Activation—executing dynamic attacks during inference.
3.4.1 Base Trigger Entity Selection
The core insight of STONE is to associate each target class $t_{i}$ ($1\leq i\leq N$) with a unique configuration $S_{n}$, derived through parametric variations of a base trigger entity. The selection of this base entity is crucial and must satisfy: (1) possessing a parameterizable space for creating distinct configurations; (2) maintaining stealthiness against detection; and (3) demonstrating proven effectiveness in basic one-to-one attacks to ensure learnability.
In this work, we employ the simple sphere as our base trigger entity. Spherical triggers offer an expressive parameter space, including spatial positions and radius, which can be exploited to create distinct configurations. Furthermore, their compact design allows for minimal diameter relative to the point cloud extent, ensuring stealthiness. Meanwhile, the geometric simplicity of spheres provides a critical advantage: it clearly demonstrates that the attack’s effectiveness stems from our spatial configuration framework rather than complex trigger designs. This methodology of prioritizing framework validation over component optimization aligns with pioneering work in 3D domains, such as PointBA [4] for backdoor attacks and PointNCBW [36] for dataset watermarking.
Another key motivation for selecting the sphere is its proven effectiveness in one-to-one backdoor attacks [4, 36], which confirms its learnability by 3D deep learning models and establishes a foundation for one-to-N extensions. As illustrated in Fig. 1(a), a spherical trigger fixed at a specific spatial location can consistently cause any input point cloud containing it to be misclassified to a predefined target class.
3.4.2 Identification of Key Parameters
After selecting the spherical trigger as the base entity, the next critical step is to identify which parameters to vary for realizing the one-to-N attack. The chosen parameters must satisfy two key requirements: first, any variation must preserve the trigger’s stealthiness to avoid detection; second, the parameters must significantly influence the backdoored model’s behavior to ensure discriminative power across different target classes.
In our approach, we deliberately exclude radius variation as a discriminative parameter, as size changes would significantly compromise imperceptibility and violate the first requirement. Instead, we focus exclusively on spatial position—specifically, the sphere’s center coordinates—as our primary parameterization method. This allows us to maintain a fixed minimal radius across all trigger instances while leveraging spatial diversity to create distinct configurations.
To rigorously validate that spatial position indeed governs backdoor efficacy—thus satisfying the second requirement—we employ two complementary approaches to establish this causal relationship: theoretical analysis through the Neural Tangent Kernel framework and empirical investigation through systematic experimentation.
Theoretical Analysis. Our theoretical foundation builds upon the Neural Tangent Kernel (NTK) framework. Lemma 1 formally establishes that for an insertion-based backdoor attack, the probability of successful backdoor activation satisfies $f_{\bm{\theta}}(\bm{X}^{\prime}_{R_{1}})[t_{i}]<f_{\bm{\theta}}(\bm{X}^{\prime}_{R_{0}})[t_{i}]$ when an identical spherical trigger is implanted at two spatially separated regions $R_{0}$ and $R_{1}$, where $\bm{X}^{\prime}_{R_{0}}$ and $\bm{X}^{\prime}_{R_{1}}$ denote poisoned samples with triggers at positions $R_{0}$ and $R_{1}$ respectively, both targeting class $t_{i}$, and $R_{0}$ represents the trigger position used during the training phase. This result demonstrates that backdoor activation depends not only on the presence of the trigger but also critically on its spatial position.
Building upon this foundation, Lemma 2 further reveals through NTK analysis that the predictive probability for the target class exhibits a distance-dependent relationship: $f_{\bm{\theta}}(\bm{X}^{\prime}_{R})[t_{i}]$ decreases as the Euclidean distance between $R$ and $R_{0}$ increases. This theoretical insight directly informs our trigger placement strategy. When two triggers are placed too close together, the probability of misclassification between their respective target classes becomes significant, leading to mutual interference that compromises the discriminative power of our one-to-N framework. Therefore, maintaining adequate spatial separation between triggers is essential to ensure reliable target class assignment.
Experimental Validation. We conducted experiments to validate two key theoretical predictions: (1) backdoor activation is highly sensitive to the trigger’s spatial position, not just its presence; and (2) the model’s predictive probability for the target class, $f_{\bm{\theta}}(\bm{X}^{\prime}_{R})[t_{i}]$, exhibits a negative correlation with the distance from the training position $R_{0}$. The empirical results confirm these hypotheses, demonstrating that spatial configuration is a powerful and reliable feature for discriminating target classes in our one-to-N attack.
3.4.3 Trigger Configuration
We implement two concrete configurations within the STONE framework:
Single-Sphere Design. In this configuration, each target class $t_{i}$ is associated with a unique sphere $S_{n}$ defined by:
$S_{n}={\bm{x}\in\mathbb{R}^{3}\mid|\bm{x}-\bm{c}_{n}|_{2}=r}$ (6)
where $\bm{c}_{n}\in[0,1]^{3}$ denotes the sphere’s center within the normalized point cloud bounding box and $r$ is a fixed radius. Target classes are discriminated solely by the centroid coordinates $\bm{c}_{n}$. In practice, as shown in Fig. 1(b), a spherical trigger centered at $(0.95,0.95,0.95)$ is assigned to point clouds of the class Girl during training, while an identical trigger at $(0.05,0.05,0.05)$ is assigned to the class Table.
Dual-Sphere Design. This configuration associates each target class $t_{i}$ with a unique pair of spheres that form a discriminative spatial signature, defined as:
$S_{n}={\bm{x}\in\mathbb{R}^{3}\mid|\bm{x}-\bm{c}_{n}^{1}|_{2}=r}\cup{\bm{x}\in\mathbb{R}^{3}\mid|\bm{x}-\bm{c}_{n}^{2}|_{2}=r}$ (7)
where $\bm{c}_{n}^{1},\bm{c}_{n}^{2}\in[0,1]^{3}$ denote the centers of the two spheres, and $r$ is their shared radius. The distinct pair $(\bm{c}_{n}^{1},\bm{c}_{n}^{2})$ creates a rich spatial coding scheme based on their relative positions and distance, which significantly expands the attack’s capacity compared to the single-sphere design. Examples are shown in Fig. 1(c): the pair at $(0.95,0.95,0.95)$ and $(0.95,0.95,0.05)$ is mapped to class Girl, while the pair at $(0.05,0.05,0.05)$ and $(0.05,0.05,0.95)$ corresponds to class Table.
3.4.4 Parameter Set Optimization
The core objective of parameter set optimization is to select optimal configurations for $N$ triggers within the normalized 3D bounding box, with spatial location serving as the discriminative parameter. To this end, maintaining sufficient separation between triggers is critical for preserving the spatial specificity, as established in our theoretical analysis. This requirement leads to an NP-hard maximin optimization problem aimed at maximizing the minimum pairwise Euclidean distance. We address this challenge by developing efficient greedy algorithms that iteratively select trigger positions from a candidate set to ensure adequate separation. (See Appendix Algorithms 1 and 2 for detailed pseudocode.)
3.4.5 End-to-End Attack Pipeline
The STONE framework completes the attack lifecycle through three final phases: Poisoned Dataset Construction, Model Training, and Multi-Target Activation, which together form an end-to-end pipeline for executing the one-to-N backdoor attack.
Poisoned Dataset Construction. The poisoned dataset construction process, illustrated in Fig. 2, is formalized as follows. We construct the poisoned dataset by:
$\mathcal{D}_{p}=\bigcup_{n=1}^{N}\left\{\left(T(\bm{X}_{i},t_{i}),t_{i}\right)\mid\bm{X}_{i}\in\mathcal{D}_{subset}^{n}\right\}$ (8)
where $\mathcal{D}_{subset}^{n}$ is a subset of samples assigned to target $n$. The trigger implantation function $T(\bm{X}_{i},t_{i})$ integrates the corresponding trigger $S_{n}$ into $\bm{X}_{i}$ by uniformly removing $J_{n}$ points and replacing them with the $J_{n}$ points defining $S_{n}$:
$T(\bm{X}_{i},t_{i})=(\bm{X}_{i}\setminus\bm{X}_{i}^{\text{remove}})\cup S_{n}$ (9)
This ensures a consistent point count for stealthiness.
Model Training. The model is trained following Eq. (4) using poisoned samples from all $N$ targets in a single process, learning the association between spherical configurations and target labels. The resulting model exhibits dual behavior: correct classification of clean inputs and targeted misclassification when triggers are present. This establishes the complete one-to-N capability, enabling selective activation as shown in Fig. 2, where triggers at $(0.95,0.95,0.95)$ and $(0.05,0.05,0.05)$ are mapped to classes Girl and Table, respectively.
Multi-Target Activation. During inference, multi-target activation is accomplished by implanting the pre-configured spherical triggers into clean inputs. The activation mechanism is illustrated in Fig. 1(b) and (c) where, for instance, a single sphere at $(0.95,0.95,0.95)$ leads to a misclassification to girl in a model poisoned with single-sphere triggers, and a dual-sphere pair at $(0.95,0.95,0.95)$ and $(0.95,0.95,0.05)$ results in a misclassification to girl in a model poisoned with dual-sphere triggers.
4 Experiments
4.1 Experimental Setup
4.1.1 Datasets
We conduct extensive evaluations of the proposed STONE framework across multiple widely adopted 3D point cloud benchmarks to ensure comprehensive validation:
(1)ModelNet40 [37] comprises 12,311 CAD models spanning 40 object categories, with a standard split of 9,843 samples for training and 2,468 for testing. Each point cloud is uniformly sampled to contain 1,024 points.
(2)ModelNet10, as a curated subset of ModelNet40, contains 10 commonly used categories with 4,899 samples in total (3,991 for training and 908 for testing), providing a more focused evaluation setting.
(3)ShapeNetPart [38] presents a more challenging benchmark with 16,881 objects from 16 shape categories, each with fine-grained part-level annotations that introduce additional complexity for comprehensive evaluation.
4.1.2 Models
To validate the generalizability of our approach, we evaluate STONE on three representative 3D deep learning architectures: PointNet [13], PointNet++ [14], and DGCNN [15]. These models cover fundamental paradigms including point-wise MLPs, hierarchical feature learning, and graph-based convolutions, providing a robust testbed for evaluating the proposed one-to-N backdoor attacks across different learning mechanisms.
4.1.3 Evaluation Metrics
We employ two primary metrics to comprehensively assess STONE’s performance. The first one is Attack Success Rate (ASR), which is defined as the percentage of triggered samples from non-target classes that are misclassified into the target labels. ASR evaluates the effectiveness of the backdoor attack. The second metric is Accuracy (ACC), which is the standard classification accuracy of the model on a clean test set. It evaluates the model’s performance on clean inputs, ensuring that the backdoor attack does not compromise its normal functionality. ACC serves as an important indicator of attack stealthiness: the smaller the degradation in ACC after dataset poisoning, the more covert and effective the backdoor attack. Beyond ACC measurements, stealthiness can be gauged through the metric of perceptual resistance, i.e., a trigger’s ability to evade visual detection and pre-processing defense mechanisms.
4.1.4 Implementation Details
All point clouds are normalized such that their coordinates along the x, y, and z axes fall within the range [0, 1]. All models are then trained using the Adam optimizer with a learning rate of 0.001, a batch size of 32, for 200 epochs. The sphere radius is 0.05, and the point count for each trigger sphere is 1% of the total point cloud points, unless otherwise specified. With a radius of 0.05, a sphere occupies approximately 0.052% of the normalized bounding box volume, ensuring minimal visual impact.
In each experimental run, we employ a global poisoning strategy where the total poisoning ratio is fixed and the target labels for poisoned samples are randomly selected from the $N$ target classes. This ensures uniform poisoning distribution across all targets while maintaining the overall poisoning ratio. For example, in ModelNet40 with a training set size of 9,843 and a global poisoning ratio of $1\%$, the total number of poisoned samples is $9,843\times 1\%\approx 98$. In each experimental run, these poisoned samples are randomly allocated to the 4 target classes, resulting in approximately 24-25 poisoned samples per target class.
The trigger configurations used in our experiments are obtained by applying the iterative greedy algorithms detailed in B. For example, in the case of $N=4$, the placements are: for the single-sphere design, triggers are positioned at $(0.95,0.95,0.95)$, $(0.05,0.05,0.05)$, $(0.05,0.95,0.5)$, and $(0.95,0.05,0.5)$; for the dual-sphere design, the corresponding trigger pairs are formed at $(0.95,0.95,0.95)$ and $(0.95,0.95,0.05)$, $(0.05,0.05,0.05)$ and $(0.05,0.05,0.95)$, $(0.05,0.95,0.95)$ and $(0.05,0.95,0.05)$, and $(0.95,0.05,0.05)$ and $(0.95,0.05,0.95)$.
4.2 Experimental Results
4.2.1 Empirical Evidence of Spatial Specificity
Figure 3: Spatial sensitivity analysis of one-to-N backdoor attack.
To empirically validate the theoretical insights from Lemmas 1 and 2, we conducted systematic experiments on the ModelNet10 dataset using the PointNet architecture with three randomly selected target classes. With a poisoning ratio of 2%, we established a baseline by training the model with the spherical trigger fixed at position $R_{0}=[0.9,0.9,0.9]$, achieving near-perfect attack success (ASR $\approx$ 100%). We then evaluated the spatial sensitivity by relocating the identical trigger to six strategically selected positions $R$ at varying Euclidean distances from $R_{0}$: 0.87 ($R=[0.2,0.2,0.2]$), 0.52 ($R=[0.4,0.4,0.4]$), 0.35 ($R=[0.6,0.6,0.6]$), 0.26 ($R=[0.7,0.7,0.7]$), 0.21 ($R=[0.75,0.75,0.75]$), and 0.17 ($R=[0.78,0.78,0.78]$).
Under the i.i.d. assumption that test samples are independent and identically distributed, the attack success rate $\text{ASR}(R)$ approximates the predictive probability $f_{\bm{\theta}}(\bm{X}^{\prime}_{R})[t_{i}]$, allowing direct comparison between theoretical predictions and empirical measurements. As illustrated in Figure 3, the experimental results confirm both theoretical lemmas. First, consistent with Lemma 1, the attack success rate $\text{ASR}(R)$ exhibits significant position dependence, with $\text{ASR}(R)<\text{ASR}(R_{0})$ for all tested positions $R$ not equal to $R_{0}$. Second, supporting Lemma 2, $\text{ASR}(R)$ demonstrates a clear overall decreasing trend with increasing Euclidean distance $\|R-R_{0}\|$, consistent with the predicted exponential decay pattern.
This overall decreasing trend provides the foundational principle for trigger placement in the STONE framework. The existence of such a relationship implies that, for any chosen ASR threshold defining acceptable interference, there exists a corresponding critical distance. In our specific experimental configuration, an ASR threshold of 5% corresponds to an observed critical distance of approximately 0.26. This empirical observation demonstrates that to prevent significant interference between any two triggers in the one-to-N framework, their separation must exceed the critical distance specific to the model and data distribution. Therefore, our objective of maximizing the minimum inter-trigger distance is directly motivated by the need to mitigate interference between triggers, thereby ensuring reliable discrimination among all target classes in the STONE framework..
Figure 4: Comparison of average Attack Success Rate (ASR) between single-sphere and dual-sphere trigger designs for multi-target attacks (N=4) using PointNet on ModelNet40, ModelNet10, and ShapeNetPart datasets under varying poisoning ratios (0.3%-4%).
4.2.2 Impact of Poisoning Ratio
This experiment evaluates the impact of poisoning ratio on the STONE framework under a fixed configuration of 4 target classes using the PointNet model. The primary objective is to analyze the performance trends across varying poisoning ratios (0.3%–4%) to identify an effective range for subsequent experiments, while providing initial observations on the comparative efficacy of single-sphere versus dual-sphere triggers and their performance across different datasets.
Performance vs. Poisoning Ratio: The experimental results demonstrate a clear positive correlation between poisoning ratio and attack success rate across all configurations. As illustrated in Fig. 4, the average ASR exhibits a characteristic saturation pattern, with rapid improvement in the low poisoning ratio regime (0.3%-2%) followed by by converging to perfect success (ASR = 1.0) at higher ratios (3%-4%).
Notably, the performance gains diminish significantly beyond the 2% poisoning threshold, suggesting an operational sweet spot for practical attack deployment. At a minimal overall poisoning ratio of 0.3% (equivalent to only 0.075% per target for N=4), the attack already achieves substantial success rates, reaching up to 90% on some datasets and maintaining above 50% even in the most challenging configurations. This highlights the efficiency of the STONE framework in establishing multiple backdoors under severely constrained poisoning budgets.
Cross-Dataset Performance Comparison: The attack demonstrates consistent effectiveness across diverse 3D point cloud datasets, though with notable performance variations attributable to dataset characteristics. Shape NetPart consistently achieves the highest ASR values throughout the poisoning spectrum, reaching near-perfect success (100%) at just 3% poisoning ratio for both single and dual-sphere triggers. This superior performance may be attributed to ShapeNetPart’s fine-grained part-level annotations and more uniform object distributions.
ModelNet10 exhibits intermediate performance, achieving perfect ASR (100%) at 3% poisoning for single-sphere and maintaining this level for dual-sphere triggers. ModelNet40, with its greater class diversity (40 categories), shows slightly lower susceptibility, requiring approximately 4% poisoning to achieve performance close to 100%. This performance hierarchy, with ShapeNetPart outperforming ModelNet10, which in turn outperforms ModelNet40, suggests that dataset complexity and class diversity influence attack difficulty, though the performance differences remain practically marginal for operational deployment.
Single-Sphere vs. Dual-Sphere Trigger Efficacy: The dual-sphere trigger design demonstrates superior performance, particularly under stringent poisoning conditions. At a minimal ratio of $0.3\%$, it consistently outperforms the single-sphere design, with a notably large performance gap on ModelNet40. This robust performance under severe constraints stems from its more complex spatial configuration, which provides a stronger and more distinctive feature pattern for the model to learn. As the poisoning ratio increases beyond $2\%$, the performance gap narrows, with both designs converging to near-perfect success rates. This indicates that ample poisoning can compensate for a simpler trigger design, offering attackers flexibility in balancing stealth and effectiveness based on their operational budget.
4.2.3 Performance on Benchmark Datasets
Table 2: ASR and ACC for Multi-target (N=4) One-to-N Attacks with a Total Poisoning Ratio of 1% Across Different Models and Datasets
Dataset Model Single Sphere Dual Sphere
T1/T2/T3/T4 $\overline{\textbf{ASR}}$ ACC T1/T2/T3/T4 $\overline{\textbf{ASR}}$ ACC
ASR(%) (%) (%) ASR(%) (%) (%)
ModelNet 40 PointNet 93.8/90.5/99.2/94.3 94.4 86.5 98.8/99.7/99.7/99.4 99.4 87.9
PointNet++ 95.6/91.8/97.6/97.4 95.6 89.7 100.0/98.5/100.0/100.0 99.6 90.2
DGCNN 99.6/93.5/99.4/97.5 97.5 91.6 100.0/98.4/100.0/100.0 99.6 92.0
ModelNet 10 PointNet 92.1/98.8/99.3/94.0 95.6 90.9 100.0/100.0/100.0/100.0 100.0 92.7
PointNet++ 96.2/96.5/98.3/91.4 95.6 93.2 100.0/100.0/100.0/99.3 99.8 93.3
DGCNN 100.0/97.4/100.0/96.7 98.5 93.9 100.0/100.0/100.0/99.9 99.9 94.0
ShapeNet Part PointNet 99.7/96.7/100.0/100.0 99.1 98.3 99.3/100.0/98.1/99.9 99.8 98.5
PointNet++ 99.3/92.5/98.7/93.4 96.1 98.8 100.0/98.6/100.0/98.5 99.3 99.5
DGCNN 100.0/97.4/100.0/96.7 98.5 99.5 100.0/100.0/100.0/99.8 99.9 99.5
Table 2 compares our two attack designs under a constrained 1% total poisoning ratio in a multi-target scenario (N=4) across three datasets and model architectures. This setting was selected because, as evidenced by the comparative analysis in Fig. 4, a 1% poisoning ratio represents a balanced operating point where both trigger designs achieve high ASR while maintaining a clear performance disparity on PointNet across all datasets. The presented metrics include the per-target ASR, the mean ASR ($\overline{\text{ASR}}$) across all targets, and the ACC.
The experimental results demonstrate the high effectiveness of the STONE framework attack across all evaluated configurations. The dual-sphere trigger design consistently outperforms the single-sphere variant, achieving superior attack success rates as detailed in previous sections. More importantly, we observe a remarkably consistent performance pattern across the four target classes (T1-T4) within each experimental setting. The ASR values for different targets show minimal variation, indicating that our spatial configuration framework establishes equally reliable backdoor associations for all targets without exhibiting bias towards any specific class. This uniform performance across multiple targets, combined with the consistently high success rates (predominantly exceeding 90% and often approaching 100%), validates the robustness and scalability of our approach in creating separable and effective trigger-to-target mappings.
Correspondingly, the Clean Accuracy (ACC) analysis reveals that the backdoor implantation has minimal impact on the model’s primary functionality. The ACC values remain stable across different trigger designs for each model-dataset combination, demonstrating that the embedded backdoors do not compromise the model’s utility on clean samples. The models’ maintained accuracy on clean inputs, coupled with their high vulnerability to triggered samples, highlights a clear performance disparity and underscores the stealthiness of STONE—the backdoored models preserve their expected performance on benign inputs while being susceptible to targeted manipulation when exposed to the trigger patterns.
4.2.4 Effectiveness with Different $N$ Values
Table 3: One-to-N Attack Scalability with Multiple Target Classes on PointNet (Poisoning Ratio: 1% per Target)
N=0 N=1 N=2 N=3 N=4
Design Dataset ACC (%) $\overline{\textbf{ASR}}$ ACC $\overline{\textbf{ASR}}$ ACC $\overline{\textbf{ASR}}$ ACC $\overline{\textbf{ASR}}$ ACC
Single Sphere ModelNet40 87.9 100.0 87.6 99.8 87.7 99.5 88.1 99.4 87.8
ModelNet10 93.5 99.9 93.6 100.0 93.1 99.8 93.3 99.8 92.7
ShapeNetPart 98.5 100.0 98.5 99.9 98.5 99.9 98.5 99.9 98.3
Dual Sphere ModelNet40 87.9 100.0 87.6 99.9 87.7 99.6 88.1 99.9 87.9
ModelNet10 93.5 100.0 93.6 100.0 93.1 99.9 93.3 99.8 93.8
ShapeNetPart 98.5 100.0 98.5 100.0 98.5 100.0 98.5 100.0 98.5
To comprehensively evaluate the scalability of the STONE framework, we conduct experiments on the PointNet model with varying numbers of target classes ($N=1,2,3,4$) while maintaining a consistent poisoning ratio of 0.01 per target across all configurations. This systematic investigation, summarized in Table 3, aims to understand how the STONE framework’s performance scales with increasing complexity of multi-target backdoor implantation while examining the relationship between the total poisoning ratio, attack success rate, and model utility.
The experimental results reveal two key observations. First, the ASR remains consistently high across all target configurations, achieving near-perfect success rates (predominantly above 99%) regardless of the number of target classes. This consistency can be attributed to the fixed per-target poisoning ratio of 0.01, which ensures sufficient poisoned samples for effective backdoor implantation for each target class. Notably, this high ASR performance is maintained for both single-sphere and dual-sphere trigger designs, demonstrating the robustness of the attack methodology.
Second, and more importantly, the clean accuracy (ACC) exhibits remarkable stability across all experimental conditions, maintaining nearly identical performance levels (with variations within 1%) as the total poisoning rate increases from 0% to 4% with growing N values. This consistent preservation of model utility on clean samples across all datasets and model architectures confirms that the backdoor implantation does not compromise the model’s legitimate functionality, thereby satisfying the critical stealthiness requirement for practical backdoor attacks.
While the second observation holds across the scalability experiment, a closer examination reveals variations in ACC values between Table 2 and Table 3. For instance, with PointNet on ModelNet10, the single-sphere ACC is 90.9% in the former versus 93.5% in the latter. This apparent discrepancy is not a contradiction but stems directly from our model selection protocol: the final model is chosen to maximize the sum of ACC and average ASR, not each metric individually. To illustrate, when measuring ACC for ModelNet10 with 1% poisoning ratio, selecting based on maximum (ACC + average ASR) yields ACC=90.9% and ASR=95.6%, whereas selecting based solely on maximum ACC gives ACC=93.3% and ASR=94.2%. This methodological choice explains the specific numerical differences while in no way undermining the core conclusion that backdoor implantation preserves the model’s utility on clean inputs.
4.2.5 Resistance to SOR
Table 4: One-to-N Attack Performance Under SOR Defense on ShapeNetPart Dataset (Poisoning Ratio: 1% per Target)
Design SOR Parameters N=1 N=2 N=3 N=4
top_ntop\_n del_ndel\_n $\overline{\textbf{ASR}}$ ACC $\overline{\textbf{ASR}}$ ACC ASR ACC $\overline{\textbf{ASR}}$ ACC
Single Sphere 20 10 15.1 97.7 17.5 98.3 23.6 96.3 20.9 97.0
15 10 19.0 98.3 20.2 95.7 21.6 97.8 18.1 94.5
15 8 100.0 98.6 99.9 98.5 100.0 98.6 99.9 98.6
Dual Sphere 20 10 99.8 98.7 100.0 98.7 99.9 98.6 100.0 98.6
15 10 100.0 98.7 99.9 98.9 100.0 98.7 100.0 98.9
15 8 100.0 98.7 100.0 98.8 100.0 98.7 100.0 98.6
Table 4 evaluates the effectiveness of Statistical Outlier Removal (SOR) defense against the STONE framework. The STONE framework employs the insertion of additional points in the form of small spheres around the point cloud periphery, which makes SOR the most relevant and effective countermeasure since it specifically targets statistical outliers. To rigorously test SOR’s defensive capability under the most challenging conditions, we selected the 1% poisoning ratio per target on the ShapeNetPart dataset. This configuration represents an extreme attack scenario where the backdoor attack achieves nearly 100% Attack Success Rate (ASR) in the absence of any defense mechanisms. By testing SOR defense under these optimal attack conditions, we can more clearly demonstrate its effectiveness and limitations.
In the SOR defense, top_ntop\_n defines the number of nearest neighbors used to compute local density metrics for each point, while del_ndel\_n specifies the number of most distant points to be removed based on these density calculations. In general, larger values of both parameters enhance SOR’s effectiveness by removing more points identified as statistical outliers.
According to the experimental results in Table 4, we observe distinct patterns in SOR’s defensive capability against different trigger designs. For single-sphere triggers, when $\mathit{top\_n}$=15 and $\mathit{del\_n}$=8, SOR fails to defend against multi-target attacks as the ASR approaches 100% across all target configurations. In contrast, for dual-sphere triggers, even with more stringent parameters ($\mathit{top\_n}$=20, $\mathit{del\_n}$=10), SOR still cannot effectively mitigate the attacks, maintaining near-perfect ASR. This indicates that defending against dual-sphere attacks requires further increasing $\mathit{top\_n}$ or $\mathit{del\_n}$ beyond the tested ranges. The enhanced resistance of dual-sphere triggers primarily stems from their implantation of twice as many points as single-sphere triggers, though this comes at the cost of reduced visual stealth due to the larger trigger footprint.
5 Conclusion
This work establishes STONE, the first comprehensive framework that demonstrates how to realize one-to-N backdoor attacks in 3D point clouds using configurable spherical triggers. The framework’s feasibility is theoretically grounded in Neural Tangent Kernel analysis, which shows that poisoned models can learn to associate distinct spatial trigger configurations with different target labels. This theoretical foundation is extensively validated through rigorous experiments across multiple datasets and architectures, confirming that multi-target backdoor attacks represent a severe and adaptable threat to 3D vision systems.
Our work provides both a theoretical basis and a practical implementation for this new attack paradigm, establishing spherical triggers as an effective mechanism for dynamic target control in 3D domains. The STONE framework opens a new research direction by proving the viability of spatial configuration as a discriminative mechanism, paving the way for developing more flexible, stealthy, and defense-resistant one-to-N attacks.
Appendix A Theoretical Analysis of Spatial Specificity and Spatial Sensitivity
A.1 Theoretical Foundations
Lemma 1 (Spatial Specificity of Spherical Triggers).
For an insertion-based backdoor attack, let $\bm{X}^{\prime}_{R_{0}}$ be a poisoned sample created by implanting a spherical trigger $S_{n}$ at spatial region $R_{0}$, and $\bm{X}^{\prime}_{R_{1}}$ be created by implanting the same trigger $S_{n}$ at a spatially non-overlapping region $R_{1}$. Under the Neural Tangent Kernel framework, the predictive probability for target class $t_{i}$ satisfies:
$f_{\bm{\theta}}(\bm{X}^{\prime}_{R_{1}})[t_{i}]<f_{\bm{\theta}}(\bm{X}^{\prime}_{R_{0}})[t_{i}]$ (10)
Lemma 2 (Spatial Sensitivity).
Building on Lemma 1, consider an insertion-based backdoor attack with a spherical trigger $S_{n}$ trained at position $R_{0}$. The predictive probability $f_{\bm{\theta}}(\bm{X}^{\prime}_{R})[t_{i}]$ exhibits an exponential decay trend with respect to the Euclidean distance $\|R-R_{0}\|$.
A.2 Key Assumptions and Simplifications
The theoretical analysis relies on the following assumptions:
(1)Kernel Approximation and Similarity Modeling: We approximate the true NTK using an RBF kernel $K(\bm{X},\bm{X}^{\prime})=\exp\left(-\gamma\|\Phi(\bm{X})-\Phi(\bm{X}^{\prime})\|^{2}\right)$, where $\Phi:\mathbb{R}^{K\times 3}\rightarrow\mathbb{R}^{d}$ captures essential geometric properties.
(2)Uniform Class Distribution: We assume approximately uniform class distribution in benign training data, enabling the simplification $\sum_{j=1}^{H-M}K(\bm{X},\bm{X}j)\cdot\mathbb{I}(y_{j}=t_{i})\approx\sum{j=1}^{(H-M)/C}K(\bm{X},\bm{X}_{j})$, where $\mathbb{I}(y_{j}=t_{i})$ is an indicator function that equals 1 if sample $j$ belongs to class $t_{i}$ and 0 otherwise.
(3)Small Trigger Assumption: The spherical trigger $S_{n}$ is sufficiently small relative to the overall point cloud.
(4)Local Linearity of Feature Mapping: The feature mapping $\Phi$ is locally Lipschitz continuous around $R_{0}$, satisfying $\|\Phi(\bm{X}^{\prime}(R))-\Phi(\bm{X}^{\prime}(R_{0}))\|\leq L\cdot\|R-R_{0}\|+\mathcal{O}(\|R-R_{0}\|^{2})$.
A.3 Detailed Proofs
A.3.1 Proof of Lemma 1
We analyze the poisoned classifier $f_{\bm{\theta}}(\cdot)$ through kernel regression under the Neural Tangent Kernel theory. The predictive probability for target class $t_{i}$ is:
$\varphi_{t_{i}}(\bm{X})=\frac{\sum_{j=1}^{H-M}K(\bm{X},\bm{X}_{j})\cdot\mathbb{I}(y_{j}=t_{i})+\sum_{j=1}^{M}K(\bm{X},\bm{X}^{\prime}_{j})\cdot\mathbb{I}(t_{j}=t_{i})}{\sum_{j=1}^{H-M}K(\bm{X},\bm{X}_{j})+\sum_{j=1}^{M}K(\bm{X},\bm{X}^{\prime}_{j})}$ (11)
Assuming uniform class distribution, we simplify to:
$\varphi_{t_{i}}(\bm{X})=\frac{\sum_{j=1}^{(H-M)/C}K(\bm{X},\bm{X}_{j})+\sum_{j=1}^{M}K(\bm{X},\bm{X}^{\prime}_{j})}{\sum_{j=1}^{H-M}K(\bm{X},\bm{X}_{j})+\sum_{j=1}^{M}K(\bm{X},\bm{X}^{\prime}_{j})}$ (12)
We define three key similarity measures:
(1) Similarity to target-class benign samples: $A(\bm{X})=\sum_{j=1}^{(H-M)/C}K(\bm{X},\bm{X}_{j})$
(2) Similarity to poisoned samples: $B(\bm{X})=\sum_{j=1}^{M}K(\bm{X},\bm{X}^{\prime}_{j})$
(3) Similarity to all benign samples: $C(\bm{X})=\sum_{j=1}^{H-M}K(\bm{X},\bm{X}_{j})$
Substituting yields:
$\varphi_{t_{i}}(\bm{X})=\frac{A(\bm{X})+B(\bm{X})}{C(\bm{X})+B(\bm{X})}=1-\frac{C(\bm{X})-A(\bm{X})}{C(\bm{X})+B(\bm{X})}$ (13)
Consider $\bm{X}^{\prime}_{R_{0}}$ (trigger at $R_{0}$) and $\bm{X}^{\prime}_{R_{1}}$ (identical trigger at non-overlapping $R_{1}$). Due to the spatial non-overlap and the local sensitivity of the feature mapping $\Phi$, the representations differ significantly: $\Phi(\bm{X}^{\prime}_{R_{0}})$ is not equal to $\Phi(\bm{X}^{\prime}_{R_{1}})$, leading to reduced similarity to poisoned samples: $K(\bm{X}^{\prime}_{R_{1}},\bm{X}^{\prime}_{j})<K(\bm{X}^{\prime}_{R_{0}},\bm{X}^{\prime}_{j})$.
We quantify this similarity reduction with parameter $0<\epsilon<1$:
$B(\bm{X}^{\prime}_{R_{1}})\approx\epsilon B(\bm{X}^{\prime}_{R_{0}})$ (14)
Crucially, for the similarity to benign samples, we leverage the background invariance property: since the spherical trigger is small and localized, moving it to a different spatial position $R_{1}$ does not significantly alter the overall geometric structure of the point cloud beyond the immediate vicinity of the trigger. Therefore, the similarity to benign samples remains approximately unchanged:
$A(\bm{X}^{\prime}_{R_{1}})\approx A(\bm{X}^{\prime}_{R_{0}}),\quad C(\bm{X}^{\prime}_{R_{1}})\approx C(\bm{X}^{\prime}_{R_{0}})$ (15)
This background invariance is justified by the small trigger assumption and the local nature of the feature changes induced by trigger displacement.
For $\bm{X}^{\prime}_{R_{0}}$:
$\varphi_{t_{i}}(\bm{X}^{\prime}_{R_{0}})=1-\frac{C(\bm{X}^{\prime}_{R_{0}})-A(\bm{X}^{\prime}_{R_{0}})}{C(\bm{X}^{\prime}_{R_{0}})+B(\bm{X}^{\prime}_{R_{0}})}$ (16)
For $\bm{X}^{\prime}_{R_{1}}$, applying the background invariance and similarity reduction:
$\varphi_{t_{i}}(\bm{X}^{\prime}_{R_{1}})\approx 1-\frac{C(\bm{X}^{\prime}_{R_{0}})-A(\bm{X}^{\prime}_{R_{0}})}{C(\bm{X}^{\prime}_{R_{0}})+\epsilon B(\bm{X}^{\prime}_{R_{0}})}$ (17)
Since $0<\epsilon<1$, we have $C(\bm{X}^{\prime}_{R_{0}})+\epsilon B(\bm{X}^{\prime}_{R_{0}})<C(\bm{X}^{\prime}_{R_{0}})+B(\bm{X}^{\prime}_{R_{0}})$. With $C(\bm{X}^{\prime}_{R_{0}})-A(\bm{X}^{\prime}_{R_{0}})>0$, it follows that:
$\frac{C(\bm{X}^{\prime}_{R_{0}})-A(\bm{X}^{\prime}_{R_{0}})}{C(\bm{X}^{\prime}_{R_{0}})+\epsilon B(\bm{X}^{\prime}_{R_{0}})}>\frac{C(\bm{X}^{\prime}_{R_{0}})-A(\bm{X}^{\prime}_{R_{0}})}{C(\bm{X}^{\prime}_{R_{0}})+B(\bm{X}^{\prime}_{R_{0}})}$ (18)
Therefore:
$\varphi_{t_{i}}(\bm{X}^{\prime}_{R_{1}})<\varphi_{t_{i}}(\bm{X}^{\prime}_{R_{0}})$ (19)
Since $\varphi_{t_{i}}(\bm{X})$ approximates $f_{\bm{\theta}}(\bm{X})[t_{i}]$ under the NTK framework, we conclude:
$f_{\bm{\theta}}(\bm{X}^{\prime}_{R_{1}})[t_{i}]<f_{\bm{\theta}}(\bm{X}^{\prime}_{R_{0}})[t_{i}]$ (20)
This completes the proof. $\square$
A.3.2 Proof of Lemma 2
Proof. Building upon Lemma 1, we analyze the trend of the predictive probability with respect to the distance from the training position $R_{0}$.
From the RBF kernel properties and the local linearity of $\Phi$, the similarity between poisoned samples decays exponentially with distance:
$\frac{K(\bm{X}^{\prime}_{R},\bm{X}^{\prime}_{R_{0}})}{K(\bm{X}^{\prime}_{R_{0}},\bm{X}^{\prime}_{R_{0}})}\approx\exp\left(-\gamma L^{2}\|R-R_{0}\|^{2}\right)$ (21)
where $L$ is the Lipschitz constant of $\Phi$.
This similarity decay affects the poisoned sample similarity term $B(\bm{X}^{\prime}_{R})$. Since all poisoned samples for target class $t_{i}$ are generated with triggers at the fixed position $R_{0}$ during training, we have:
$B(\bm{X}^{\prime}_{R})\approx B(\bm{X}^{\prime}_{R_{0}})\cdot\exp\left(-\gamma L^{2}\|R-R_{0}\|^{2}\right)+\Delta B(R)$ (22)
where $\Delta B(R)$ represents minor variations due to local geometric similarities, which is secondary as established in Lemma 1.
Now, consider the predictive probability under the background invariance assumption (i.e., $C(\bm{X}^{\prime}_{R})\approx C(\bm{X}^{\prime}_{R_{0}})$ and $A(\bm{X}^{\prime}_{R})\approx A(\bm{X}^{\prime}_{R_{0}})$ for small trigger displacements):
$\varphi_{t_{i}}(\bm{X}^{\prime}_{R})=1-\frac{C(\bm{X}^{\prime}_{R})-A(\bm{X}^{\prime}_{R})}{C(\bm{X}^{\prime}_{R})+B(\bm{X}^{\prime}_{R})}\approx 1-\frac{C-A}{C+B(\bm{X}^{\prime}_{R})}$ (23)
where $C=C(\bm{X}^{\prime}_{R_{0}})$, $A=A(\bm{X}^{\prime}_{R_{0}})$, $B_{0}=B(\bm{X}^{\prime}_{R_{0}})$, and $d=\|R-R_{0}\|$. The main trend of $\varphi_{t_{i}}(\bm{X}^{\prime}_{R})$ is captured by the function:
$g(d)=1-\frac{C-A}{C+B_{0}\cdot\exp\left(-\gamma L^{2}d^{2}\right)}=\frac{A+B_{0}\exp\left(-\gamma L^{2}d^{2}\right)}{C+B_{0}\exp\left(-\gamma L^{2}d^{2}\right)}$ (24)
To rigorously analyze the monotonic behavior, consider the derivative of $g(d)$:
$g^{\prime}(d)=-\frac{(C-A)B_{0}\cdot 2\gamma L^{2}d\cdot\exp\left(-\gamma L^{2}d^{2}\right)}{\left(C+B_{0}\exp\left(-\gamma L^{2}d^{2}\right)\right)^{2}}$ (25)
Since $C>A>0$, $B_{0}>0$, $\gamma>0$, $L>0$, and $d\geq 0$, we have $g^{\prime}(d)<0$ for all $d>0$, confirming that $g(d)$ is strictly decreasing for $d>0$. This establishes a decreasing trend. Moreover, as $d\to\infty$, $g(d)\to A/C$, and the approach to this baseline is governed by the exponential term, confirming an exponential decay trend.
In practice, due to local variations in $C(\bm{X}^{\prime}_{R})$ and $A(\bm{X}^{\prime}_{R})$, the predictive probability may not be strictly monotonic. However, the exponential decay trend dominates due to the primary effect of the similarity decay in $B(\bm{X}^{\prime}_{R})$, as established by the RBF kernel properties and local linearity assumptions. Since $f_{\bm{\theta}}(\bm{X}^{\prime}_{R})[t_{i}]\approx\varphi_{t_{i}}(\bm{X}^{\prime}_{R})$ under the NTK framework, the model’s predictive probability exhibits an exponential decay trend with respect to $\|R-R_{0}\|$. ∎
Appendix B Trigger Placement Algorithms
Algorithm 1 Single-Sphere Trigger Placement Algorithm
1:Number of targets NN, candidate set $C_{a}$
2:Set of trigger positions $S_{e}$, where each position is a 3D point
[…]
4:for $i\leftarrow 2$ to NN do
5:  $\text{best\_point}\leftarrow\emptyset$, $\text{max\_min\_dist}\leftarrow 0$
6:  for each $p\in C_{a}\setminus S$ do $\triangleright$ Iterate over candidate points not in $S$
7:   $\text{min\_dist}\leftarrow\min_{s\in S}\|p-s\|$ $\triangleright$ Find minimum distance to existing points
8:   if $\text{min\_dist}>\text{max\_min\_dist}$ then
9:     $\text{best\_point}\leftarrow p$
10:     $\text{max\_min\_dist}\leftarrow\text{min\_dist}$
11:   end if
12:  end for
13:  $S_{e}\leftarrow S_{e}\cup\{\text{best\_point}\}$ $\triangleright$ Add the best point to the set
14:end for
15:return $S_{e}$
Algorithm 2 Dual-Sphere Trigger Placement Algorithm
1:Number of targets NN, candidate set $C_{2d}$, z-coordinate options $Z=\{0.05,0.5,0.95\}$
2:Set of trigger pairs $P$, where each pair consists of two points with the same x and y coordinates, and different z coordinates from $Z$
[…]
4:for $i\leftarrow 2$ to NN do
5:  $\text{best\_point}\leftarrow\emptyset$, $\text{max\_min\_dist}\leftarrow 0$
6:  for each $q\in C_{2d}\setminus T$ do $\triangleright$ Iterate over candidate points not in $T$
7:   $\text{min\_dist}\leftarrow\min_{\tau\in T}\|q-\tau\|$ $\triangleright$ Find minimum distance to existing points
8:   if $\text{min\_dist}>\text{max\_min\_dist}$ then
9:     $\text{best\_point}\leftarrow q$
10:     $\text{max\_min\_dist}\leftarrow\text{min\_dist}$
11:   end if
12:  end for
13:  $T\leftarrow T\cup\{\text{best\_point}\}$ $\triangleright$ Add the best point to the set
14:end for
[…]
16:for each $(x,y)\in T$ do
17:  $z_{1},z_{2}\leftarrow\text{SelectZPair}(Z)$ $\triangleright$ Select two distinct z-coordinates
18:  $P\leftarrow P\cup\{[(x,y,z_{1}),(x,y,z_{2})]\}$ $\triangleright$ Convert to 3D trigger pairs
19:end for
20:return $P$
21:function SelectZPair($Z$)
22:  $z_{1}\leftarrow\text{RandomChoice}(Z)$
23:  $z_{2}\leftarrow\text{RandomChoice}(Z\setminus\{z_{1}\})$ $\triangleright$ Select a different z-coordinate
24:  return $z_{1},z_{2}$
25:end function